### v1

# Generating Driving Simulations via Conversation

Abstract
Cyber-physical systems like autonomous vehicles are tested in simulation before deployment, using domain-specific programs for scenario specification. To aid the testing of autonomous vehicles in simulation, we design a natural language interface, using an instruction-following large language model, to assist a non-coding domain expert in synthesising the desired scenarios and vehicle behaviours. We show that using it to convert utterances to the symbolic program is feasible, despite the very small training dataset. Human experiments show that dialogue is critical to successful simulation generation, leading to a 4.5 times higher success rate than a generation without engaging in extended conversation.
I Introduction
Testing autonomous vehicles exclusively on public roads, especially in near-crash scenarios, is not feasible [9]. Iterative development and testing in simulation is essential [18], especially in the early stage of the development process. Simulators like CARLA [4] are used for the research of autonomous vehicles in the iterative development process. Using CARLA, engineers can generate driving scenarios using Scenic [5]: probabilistic scenario description language such as Scenic specify a distribution over desired driving scenarios, and then sample many simulation instances for evaluation. In this way, challenging driving scenarios can be generated [11, 12], which can then be used for evaluation of control algorithms for autonomous vehicles without a physical system, in a safe and cost-effective manner.
Writing programs in Scenic to specify driving scenarios or more generally programming in domain-specific languages is challenging with steep learning curves. Domain experts with deep knowledge of edge cases and desired driving behaviours may lack proficiency in Scenic or other specification languages, or even when they do have such proficiency, they may still benefit from some assistance in their workflow. Natural language interfaces (dialogue systems), in which engineers converse with a chatbot that knows how to program in Scenic, is a useful tool for facilitating this. In this project, we aim to develop such a natural language interface, and by doing so increase engineer access to simulation-driven testing, by allowing them to interactively synthesise scenarios using natural language dialogue.
This is a challenging task for two reasons. Firstly due to data scarcity (only 32 examples of Scenic programs with natural language description are available online), approaches in developing a dialogue system using instruction-following large language models (IFLLM) [21] are not directly applicable. Secondly, the scenario description provided by the user may be underspecified: missing some details about the desired driving scenario in the description that is only exposed through interaction with the simulator. Fig. 1 illustrates this phenomenon. When a user utters “an ego vehicle overtakes the two adversary vehicles”, the dialogue system is able to synthesize the simulation instance satisfying the description, but the description itself does not provide all information about the situation, which in this case is the manner of overtaking (could be one by one, or both at the same time). The user only becomes aware about the necessity of providing such details by observing simulation instances and providing feedback to the system, which is used to update the program and in turn generate simulation instances.
To this end, this project makes the following contributions: 1) we create a dataset of English description-Scenic program pairs for a variety of driving scenarios; 2) we investigate how IFLLM can be used to synthesize Scenic programs from natural language expressions in an embodied conversation where the user converses with a chatbot and reacts to driving simulation instances; and 3) we conduct human trials to evaluate the dialogue system, and in particular the value of having multiple dialogue turns. Our results show that the dialogue system can be a valuable tool to interactively create driving scenarios.
II The Dialogue System
This section describes the dialogue system for generating driving simulations via conversations. Fig. 2 gives an overview of the system. It operates in two modes: code generation (§II-A); and conversation (§II-B), in which the user responds to simulation instances as depicted in Fig. 1.
II-A Code Generation
To generate Scenic programs from natural language descriptions, the dialogue system uses retrieval-augmented generation (RAG) [10] with in-context learning [2] to construct the prompt.
To generate a program from description , we constructed a dataset of size description-program pairs, via manual augmentation of the few exemplars available on the web.111https://github.com/BerkeleyLearnVerify/Scenic/tree/2.x/examples Each program is standardised in its coding style, with five distinct code blocks—map and model, constants, agent’s behavior, spatial relations, and scenario specification—to minimize the variation between the exemplars. The dataset is indexed via embeddings of descriptions.222BAAI/bge-small-en-v1.5
At inference time, the embedding of the user’s description is used to find the most similar description-program pairs from using maximum inner-product search:
The description with the retrieved exemplars are then used to create the prompt (see Appendix A for prompt specification) which is then used to generate the program using an :
There is no guarantee that the program will execute. To fix this, we perform error feeding: the program is passed to the simulator to attempt to create simulation instances; and if an error occurs, the NL description, and its errors are passed back to the , to attempt to self-correct the output. If after three attempts of error feeding the generated program doesn’t execute, the user is asked to provide an alternative description.
II-B Conversation
When executes, instances of the driving scenario simulation are shown to the user, who is given the opportunity to react to these by either indicating successful generation or by providing natural language feedback that corrects or refines the results, in an attempt to provide simulations that better align with the user’s communicative intent.
Feedback is in the response to the simulations and not to (which is not observable to the user): it can feature and refer to information that is not present in the Scenic program but is present in the simulation due to the nature of random sampling. Because of this, feeding only to is suboptimal due to its missing context, which cannot be given in a principled way (e.g. conditioning on the simulation video or some domain-specific trace would lead to further data scarcity which is already problematic in this domain). To alleviate this issue, we feedback an updated description by summarizing and , using a prompt (see Appendix for prompt specification), to generate the updated description . With this description, we aim to give guidance to , minimize the overall context, and not introduce hallucinations and undesired behaviours.
The updated description is used to generate the updated program following the previously described code generation procedure.
The conversation continues up to turns. If the user is still not satisfied with simulations, the conversation is evaluated as unsuccessful.
III Experiments
We conduct experiments for code generation and human evaluation for the overall dialogue system. NVIDIA A100 GPU is used for experiments.
III-A Code Generation Evaluation
III-A1 Experimental Setup
We experiment with several open-source s: [8], [14], and [17] that used exemplars to construct that’s retrieved using RAG, or at random with and without error feeding. To cope with the small dataset, we perform leave-one-out validation [1]. We record three metrics: (a) BLEU [15, 16] to measure generation precision; (b) ROUGE-L [13] to measure generation recall; and (c) Execution (EX) to measure the percentage of the generated code that is executable, all measured with error feeding (if required). More sophisticated code generation evaluation metrics like CodeBLEU [22], to abstract away from the surface form and focus on the syntactic similarity, are not considered because such metrics require adaptation of the parser to Scenic programs. This is partially circumvented by program standardisation performed in the dataset construction.
III-A2 Results and Discussion
Table I records the results for code generation. We observe that that uses prompt constructed using RAG is better than a random selection, which most of the time results in non-executable programs. Error feeding boosts performance for that are not fine-tuned on code (, ), yet it does increase inference time, which for leads to non-termination in reasonable time. High BLEU and ROUGE-L indicate a high correspondence between prediction and reference, as expected after program standardisation. For EX, is better than and , which is as expected due to supervised fine-tuning on code. Based on code generation experiments, we chose to use for human evaluation.
| BLEU | ROUGE-L | EX | |
| 64.32 | 63.23 | 71.43 | |
| 89.66 | 85.96 | 70.48 | |
| 0.09 | 5.80 | 0.00 | |
| 0.45 | 17.29 | 0.00 | |
| 56.86 | 55.86 | 40.95 | |
| 10.44 | 16.79 | 0.00 | |
| 84.54 | 79.94 | 41.90 | |
| 87.80 | 84.03 | 37.14 | |
| 10.41 | 30.74 | 1.90 | |
| 4.46 | 23.05 | 0.00 |
III-B Human Evaluation
III-B1 Experimental Setup
We perform a human evaluation of the dialogue system. Visual stimuli used for human trials is given in Fig. 3. The user is presented with a (diagrammatic) stimulus and asked to describe it. After that, CARLA is run to produce 3 simulation instances.333The dialogue system can produce many (different) instances, but to keep the stimuli contained, we limit it to 3. After observing it, the user may respond with further NL input to what they see, or check “satisfied with the scenario produced” and move to the next stimulus. The user can have up to dialogue turns. If code generation does not produce executable code, the user is asked to paraphrase their description as if it is a fresh interaction. The stimuli are categorised into three types: bypassing (e.g. the user stimulus may yield the user description “an ego overtakes an adversary”), intersections (e.g. “an ego goes left at a 4-way intersection while the adversary approaches from behind.”), and pedestrians (e.g.“an ego goes left and yields to the pedestrian”). The experiments were conducted with 20 users that have a driving license (as a proxy for domain proficiency), with each asked to describe 5 of the 25 stimuli. In total, 100 conversations were recorded: 44 intersection, 19 bypassing, and 37 pedestrian scenarios.
III-B2 Results and Discussion
Fig. 4 outlines the results of human evaluation. As the number of dialogue turns increases so does the the percentage of successfully generated scenarios: without dialogue, only 6% of the code generations were successful, while conversations of up to four turns raise this to 28%. This holds per scenario type, with or without execution errors. This suggests that the extended conversation enables successful scenario generation. Appendix B gives examples of a successful and failed scenario generation. Error analysis shows that failed scenario generation conversations had one or more of these errors:
Execution error (62.5%): the code produced does not execute and the user is asked to paraphrase one or more times. This happens due to a missing correspondence between Scenic API and the description. E.g. the description features “T-junction”, while the Scenic API would expect 3-way intersection, and the code features junction as a primitive rather than intersection.
Pass error (82%): the code executes but does not pass as a correct simulation. It is because the user’s description diverges from descriptions in the dataset: in length (training data descriptions were 30-50 tokens while user descriptions were up to 200 tokens); in syntax (training data descriptions were single sentences while users expressed multiple sentences and fragments); and in narrative style (training data descriptions were ego-centric while users expressed other perspectives).
Summarization error (27.8%): errors arising due to summarization (§II-B) like failure to include the feedback in the description or the updated description has extra information.
IV Conclusions
We have presented a dialogue system for conversational driving scenario generation. Our results show that users can in principle generate desired driving scenarios with conversation being a crucial element in reducing scenario generation errors.
For future work, we envision several developments. Firstly, there is a gap of failed to generate scenarios even with the extended conversation. To address this, further techniques could be considered, including data augmentation [3], constrained decoding [6], prompt engineering [19], and memory-based techniques for explicit state tracking [7]. Second, the current setup does not dynamically change the controller and use the interface to test it. This should be evaluated and considered testing of such new controllers in tandem to formal verification methods [20].
Appendix A Prompt Specification
Fig. 5 gives prompt to generate code (§ II-A) and Fig. 6 gives prompt to update descriptions (§ II-B).
[INST] You are a helpful assistant that translates English descriptions to Scenic programs. Scenic is a domain-specific probabilistic programming language for creating distributions over specified scenarios. For driving scenarios, each program has the following blocks:- MAP AND MODEL: importing town assets and enabling simulator;- CONSTANTS: specifying vehicle blueprint and other constants like vehicle speed, brake intensity and safety distance;- AGENT’S BEHAVIOR: describing how individual vehicles behave in the scenario;- SPATIAL RELATIONS: outlining the type of road the scenario needs to be synthesized in (e.g. having or not having intersections)- SCENARIO SPECIFICATION: creating individual vehicles and pedestrians in the specified roads, together with constraints that are required to betrue for the full simulation as well as the termination condition. Here are examples of description and programs:[DESCRIPTION-PROGRAM PAIRS]Now, please translate the following English description to a Scenic program. Just give the program. No extra information. Description: [DESCRIPTION][/INST]
[INST] You are a helpful assistant that creates updated driving scenario descriptions. Given a driving scenario description and corresponding feedback regarding the simulation, your task is to generate an updated description that incorporates the feedback. Please ensure that the updated description accurately reflects the intended action based on the feedback received and does not introduce additional information or lose information, like the number of vehicles or pedestrians in the situation. The description will outline a specific scenario or context, while the feedback will provide information about how the described scenario could have been improved or modified. Be sure to maintain the original meaning and intent of both the initial description and the feedback. Be brief and only return the updated driving scenario description.Description: [DESCRIPTION]Feedback: [FEEDBACK][/INST]
Appendix B Conversation Examples
Table II outlines successful and failed conversation to generate driving scenario in conversation experiments (§ III-B)
| Stimuli | |||||||||||||
| Turn | Feedback | Description | Simulation | ||||||||||
| 1 | N/A |
| 2 |
| 3 |
| 4 |
| Outcome: simulation generated successfully | |||||||||||||
| Stimuli | |||||||||||||
| Turn | Feedback | Description | Simulation | ||||||||||
| 1 | N/A |
| 2 |
| 3 |
| 4 |
| Outcome: failed to generate the simulation | |||||||||||||
![[Uncaptioned image]](extracted/5923057/figures/conv/stimuli-1.png)
![[Uncaptioned image]](extracted/5923057/figures/conv/stimuli-2.png)